OpenStudy (anonymous):

1/3 + 1/5 perform the indicated operation.

OpenStudy (anonymous):

How do you make the denomenators like?

OpenStudy (anonymous):

the lcm would be the denominator. 15 is 3&5's lcm. 3*5 = 15 1*5 = 5 so the first one is 5/15 plus 5*3 = 15 1*3=3 so 3/15 . 5/15 + 3/15 = 8/15 simplify if possible.

OpenStudy (anonymous):

thanks for the help.

OpenStudy (anonymous):

no problem. maybe try to help with somje of mine.